Chlorine is widely used to purify municipal water supplies and to treat swimming pool water. Suppose that the volume of a particular sample of Cl2 gas is 8.50 L at 895 torr and 24 ∘C.

A)

At what pressure will the volume equal 7.40 L if the temperature is 56 ∘C?

Express your answer using three significant figures
Answer #1

V1 = 8.50 Liters , P1 = 895 torr , T1 = 24°C = 24 + 273 = 297 K

V2 = 7.40 Liters , P2 = ? , T2 = 56° C = 56 + 273 = 339 K

Now using ideal gas equation,

P1 * V1 / T1 = P2 * V2 / T2

P2 = P1 * V1 * T2 / T1 * V2

P2 = 895 * 8.50 * 329 / 297 * 7.40

P2 = 1138.806 torr ..... Answer

3 significant figure after decimal